OpenStudy (acespeedfighter):

Nina had 17 marbles in a bag. Her mother put a handful of marbles in the bag. When Nina counts her marbles, she discovers she now has 34 of them. Which of the following equations will help Nina solve for the number of marbles, m, that her mom put in the bag? (3 points) 17 + 34 = m 17 + m = 34 17m = 34 m − 17 = 34

OpenStudy (mariamikayla):

The answer is variant b) the mother puts a number of marbles m on the 17 already existing in the bag and now there are 34 marbles in the bag so: 17+m=34 it means that moms put m=34-17 = 17 marbles in the bag

OpenStudy (kasey3121):

u there speed?